OpenStudy (anonymous):

Change 2x + 5y = –30 to slope-intercept form and state the slope and y-intercept?

OpenStudy (koikkara):

slope intercept from is y=mx+b so, 2x + 5y = -30 minus 2x from both sides now you have 5y = -2x - 30 now to get y alone you divide everything by 5 y = -2x-30/5 you can simplify to y = -2/5x - 6 so slope is -2/5 and y int = (0,-6)

OpenStudy (anonymous):

y = (-2/5)X-6 the slope is -2/5 the y intercept is -6

OpenStudy (anonymous):

Thank you:)